# Hausdorff Dimension of Singular Vectors in Function Fields

### v1

@@ 1. Introduction
$\max_{1\leq i\leq d}|b\theta_{i}-a_{i}|<\frac{1}{T^{\frac{1}{d}}}\qquad\text{and}\qquad 0<|b|\leq T.$
@@
$\max_{1\leq i\leq d}|b\theta_{i}-a_{i}|<\frac{\varepsilon}{T^{\frac{1}{d}}}\qquad\text{and}\qquad 0<|b|\leq T.$
@@
In this paper, we shall prove a function field analogue of Theorem 1.1. We first introduce the function field setting. Let $p$ be a prime, $q$ be a power of $p$, and let $\mathcal{R}=\mathbb{F}_{q}[x]$ be the ring of polynomials over $\mathbb{F}_{q}$. Let $\mathcal{K}=\mathbb{F}_{q}(x)$ be the field of rational functions over $\mathbb{F}_{q}$. We define an absolute value on $\mathcal{R}$ by $|f|=q^{\mathrm{deg}(f)}$ and extend it to an absolute value on $\mathcal{K}$ by $\left|\frac{f}{g}\right|=q^{\mathrm{deg}(f)-\mathrm{deg}(g)}$. Then the topological completion of $\mathcal{K}$ with respect to the metric $d(f,g)=|f-g|$ is the field of Laurent series $\widetilde{\mathcal{K}}$ defined by
@@
$\|b\boldsymbol{\theta}-\mathbf{a}\|<\frac{\varepsilon}{T^{\frac{1}{d}}}\qquad\text{and}\qquad 0<|b|\leq T.$
@@
    Since $\log_{q}(1+y)=y/\log q+o(y)$ as $y\to 0$, Theorem 1.2 has the same order of magnitude as in Theorem 1.1 for all small enough $\varepsilon$.
@@
We will now discuss a dynamical interpretation of our main theorem. Let $\operatorname{SL}_{d+1}(\widetilde{\mathcal{K}})$ be the group of invertible $(d+1)\times(d+1)$ matrices over $\widetilde{\mathcal{K}}$ with $|\det(g)|=1$ and let $\operatorname{SL}_{d+1}(\mathcal{R})$ be the subgroup of $\operatorname{SL}_{d+1}(\widetilde{\mathcal{K}})$ with entries in $\mathcal{R}$. Then, the homogeneous space $\mathcal{L}_{d+1}=\operatorname{SL}_{d+1}(\widetilde{\mathcal{K}})/\operatorname{SL}_{d+1}(\mathcal{R})$ is identified with the space of unimodular lattices in $\widetilde{\mathcal{K}}^{d+1}$ under the identification
@@
There are two main reasons for the difference between Theorems 1.1 and 1.2. The first reason is about the characterization of $\varepsilon$-Dirichlet improvability in terms of the length of the shortest vector in the Farey lattice given in Definition 4.9 with respect to best approximations. This corresponds to [CC16, Corollary 4.4]. In the real case, the characterization works upto $1/2$ and $2$ multiplications, but the ultrametric nature allows us to obtain the exact equivalence as in Corollary 4.7. The second reason comes from Minkowski’s second theorem (Theorem 2.4). While Minkowski’s second theorem on $\mathbb{R}^{d}$ works with $2^{d}/d!$ and $2^{d}$ multiplications, we have the equality over $\widetilde{\mathcal{K}}^{d}$ as in Theorem 2.4. These two reasons ultimately ensure that the fractal structures we will build work well for all $10<\varepsilon<$ 1, unlike [CC16].
@@ 2.1. Basic number theory
Given a non-zero polynomial $f\in\mathcal{R}$, let $f=\alpha P_{1}^{e_{1}}P_{2}^{e_{2}}\cdots P_{t}^{e_{t}}$ be its prime decomposition, where, $\alpha\in\mathbb{F}_{q}^{*}$, $P_{i}$ are distinct monic irreducible polynomials, and $e_{i}$ are non-negative integers. It follows from [Ros13, Propositions 1.7 and 2.4] that
@@ 2.2. Geometry of numbers
[BK23, Lemma 6.2] Let $\Lambda\subseteq\widetilde{\mathcal{K}}^{d}$ be a lattice and $B(0,r)\subseteq\widetilde{\mathcal{K}}^{d}$ be the closed ball with radius $r>0$. Then
@@
$e(X)=\sup\{r>0:\exists\alpha:B(\alpha,r)\cap X=\emptyset\}$
@@
Since $e$ is $\operatorname{SL}_{d}(\mathcal{O})$ invariant, by [Ara23, Theorem 1.15], it suffices to assume that $\Lambda=\mathcal{R}x^{\log_{q}\lambda_{1}(\Lambda)}\mathbf{e}_{1}\oplus\dots\oplus\mathcal{R}x^{\log_{q}\lambda_{d}(\Lambda)}\mathbf{e}_{d}$. Define $\mathbf{v}=\sum_{i=1}^{d}x^{\log_{q}\lambda_{i}(\Lambda)-1}\mathbf{e}_{i}$. We claim that
@@
On the other hand, by [Ara23, Theorem 1.13],
$\Lambda+B\left(0,\frac{1}{q}\lambda_{d}(\Lambda)\right)=\widetilde{\mathcal{K}}^{d}$
@@
Since $e$ is $\operatorname{SL}_{d}(\mathcal{O})$ invariant, then, by [Ara23, Theorem 1.15] , it suffices to assume that $L=\mathcal{R}x^{\log_{q}\lambda_{1}(L)}\mathbf{e}_{1}\oplus\dots\oplus\mathcal{R}x^{\log_{q}\lambda_{d}(L)}\mathbf{e}_{d}$. In this case, Theorem 2.4 implies that
@@ 3. Self-similar coverings
  • for all $\sigma$-admissible sequence $(\alpha_{k})_{k\in\mathbb{N}}$, $\lim_{k\to\infty}\operatorname{diam}B(\alpha_{k})=0$;
@@ 4.1. Best approximation vectors
See [BZ19, KLP23] for the existence of a best approximation sequence of $\boldsymbol{\theta}\in\widetilde{\mathcal{K}}^{d}$. If $\mathbf{u}=(a,b)\in Q$ is a best approximation of $\boldsymbol{\theta}\in\widetilde{\mathcal{K}}^{d}$, then every element of $\mathbb{F}_{q}^{*}\mathbf{u}$ is also a best approximation. Moreover, $b$ is unique up to $\mathbb{F}_{q}^{*}$-multiples. To see this, if $\mathbf{v}=(a^{\prime},b^{\prime})\in Q$ is a best approximation of $\boldsymbol{\theta}$ with $b^{\prime}\notin\mathbb{F}_{q}^{*}b$ and $|b^{\prime}|=|b|$, we may assume that $0<|b-b^{\prime}|<|b|$. Since $A(\boldsymbol{\theta},\mathbf{u})=A(\boldsymbol{\theta},\mathbf{v})$, we have
$\|(b-b^{\prime})\boldsymbol{\theta}-(a-a^{\prime})\|\leq\max\{A(\boldsymbol{\theta},\mathbf{u}),A(\boldsymbol{\theta},\mathbf{v})\}=A(\boldsymbol{\theta},\mathbf{u}).$
@@
By Remark 4.2, since $b$ is unique up to $\mathbb{F}_{q}^{*}$-multiples, if $b^{\prime}\notin\mathbb{F}_{q}^{*}b$, then we have $A(\boldsymbol{\theta},\mathbf{u})<A(\boldsymbol{\theta},\mathbf{v})$. Thus, the right hand side of (4.1) is equal to $A(\hat{\mathbf{u}},\mathbf{v})$, that is, $A(\boldsymbol{\theta},\mathbf{v})\leq A(\hat{\mathbf{u}},\mathbf{v})$. Suppose that $b^{\prime}\in\mathbb{F}_{q}^{*}b$ and we may assume that $b^{\prime}=b$ and $A(\boldsymbol{\theta},\mathbf{u})=A(\boldsymbol{\theta},\mathbf{v})$. Since $\mathbf{v}\notin\mathbb{F}_{q}^{*}\mathbf{u}$ and $A(\boldsymbol{\theta},\mathbf{u})\leq 1$, it follows that $A(\hat{\mathbf{u}},\mathbf{v})=\|a-a^{\prime}\|\geq 1\geq A(\boldsymbol{\theta},\mathbf{v})$.
On the other hand, by Lemma 1.5
@@
Let $\boldsymbol{\theta}\in\widetilde{\mathcal{K}}^{d}\setminus\mathcal{K}^{d}$ and let $\{\mathbf{u}_{n}\}_{n\in\mathbb{N}}$ be the sequence of best approximation vectors for $\boldsymbol{\theta}$. Then, for every $i,j\in\mathbb{N}$, we have
@@ 4.2. Orthogonality and Farey lattice
where $\pi_{\mathbf{u}}(a,b)=b\hat{\mathbf{u}}-a$, where $(a,b)\in\widetilde{\mathcal{K}}^{d}\times\widetilde{\mathcal{K}}$. Note that $\det(\Lambda_{\mathbf{u}})=|\mathbf{u}|^{-1}$. We define
@@
For $\mathbf{u}\in Q$, let $\Lambda_{\mathbf{u}}^{\prime}$ be the sublattice of $\Lambda_{\mathbf{u}}$ of minimal covolume, and denote $H_{\mathbf{u}}^{\prime}=\operatorname{span}_{\widetilde{\mathcal{K}}}(\Lambda_{\mathbf{u}}^{\prime})$ and $H_{\mathbf{u}}=\pi_{\mathbf{u}}^{-1}(H_{\mathbf{u}}^{\prime})$. Let $\xi_{1},\dots,\xi_{d}$ be an orthogonal basis of $\Lambda_{\mathbf{u}}$ with $\|\xi_{i}\|=\lambda_{i}(\Lambda_{\mathbf{u}})$ for each $i=1,\dots,d$. It can be easily checked that $\Lambda_{\mathbf{u}}^{\prime}=\mathrm{span}_{\mathcal{R}}\{\xi_{1},\dots,\xi_{d-1}\}$ and $H_{\mathbf{u}}^{\prime}=\mathrm{span}_{\widetilde{\mathcal{K}}}\{\xi_{1},\dots,\xi_{d-1}\}$.
@@
For $\mathbf{u}\in Q$, we have $\lambda_{1}(\Lambda_{\mathbf{u}})=r(\mathbf{u}).$
@@
First, observe that

### v2

@@ 1. Introduction
$\max_{1\leq i\leq d}|b\theta_{i}-a_{i}|<\frac{1}{T^{\frac{1}{d}}}\qquad\text{and}\qquad 0<|b|<T.$
@@
$\max_{1\leq i\leq d}|b\theta_{i}-a_{i}|<\frac{\varepsilon}{T^{\frac{1}{d}}}\qquad\text{and}\qquad 0<|b|<T.$
@@
In this paper, we shall prove a function field analogue of Theorem 1.1. We first introduce the function field setting. Let $p$ be a prime number, $q$ be a power of $p$, and let $\mathcal{R}=\mathbb{F}_{q}[x]$ be the ring of polynomials over $\mathbb{F}_{q}$. Let $\mathcal{K}=\mathbb{F}_{q}(x)$ be the field of rational functions over $\mathbb{F}_{q}$. We define an absolute value on $\mathcal{R}$ by $|f|=q^{\mathrm{deg}(f)}$ and extend it to an absolute value on $\mathcal{K}$ by $\left|\frac{f}{g}\right|=q^{\mathrm{deg}(f)-\mathrm{deg}(g)}$. Then the topological completion of $\mathcal{K}$ with respect to the metric $d(f,g)=|f-g|$ is the field of Laurent series $\widetilde{\mathcal{K}}$ defined by
@@
$\|b\boldsymbol{\theta}-\mathbf{a}\|<\frac{\varepsilon}{T^{\frac{1}{d}}}\qquad\text{and}\qquad 0<|b|<T.$
@@
    Since $\log_{q}(1+y)=y/\log q+o(y)$ as $y\to 0$, Theorem 1.2 has the same order of magnitude as in Theorem 1.1 for all small enough $\varepsilon$. (In fact, the lower bound of Theorem 1.2 is slightly better than $\epsilon^{t}$ in Theorem 1.1.)
@@
  4. (4)
    We note that in [CC16, Section 7], Cheung and Chevallier proved that when $d=2$, the exponent $t$ in Theorem 1.1 can be taken to be greater than $1$, instead of greater than $2$. In order to do so, they defined a more refined fractal structure, and used a claim about chord length in balls in $\mathbb{R}^{2}$ [CC16, Lemma 7.7] to show that the structure they defined is indeed nested. Since the geometry in the function field setting is non-Euclidean, similar results cannot be achieved immediately and new ideas would be needed.
One may think of the case that the base field is infinite. In this case, there have been several developments in Diophantine approximation as follows: Roth theorem [Uch60] and Schmidt subspace theorem [Rat78] over base fields of characteristic zero; a disproof of Littlewood conjecture [DL63] when the base field is infinite (see also [AB07]). However, the polynomial ring over an infinite base field behaves differently from integers. In particular, the number of polynomials within a certain radius is infinite. Since our proof relies heavily on counting polynomials, we cannot handle this case.
We will now discuss a dynamical interpretation of our main theorem. Let $\operatorname{SL}_{d+1}(\widetilde{\mathcal{K}})$ be the group of invertible $(d+1)\times(d+1)$ matrices $g$ over $\widetilde{\mathcal{K}}$ with $|\det(g)|=1$ and let $\operatorname{SL}_{d+1}(\mathcal{R})$ be the subgroup of $\operatorname{SL}_{d+1}(\widetilde{\mathcal{K}})$ with entries in $\mathcal{R}$. Then, the homogeneous space $\mathcal{L}_{d+1}=\operatorname{SL}_{d+1}(\widetilde{\mathcal{K}})/\operatorname{SL}_{d+1}(\mathcal{R})$ is identified with the space of unimodular lattices in $\widetilde{\mathcal{K}}^{d+1}$ under the identification
@@
There are two main reasons for the difference between Theorems 1.1 and 1.2. The first reason is about the characterization of $\varepsilon$-Dirichlet improvability in terms of the length of the shortest vector in the Farey lattice given in Definition 4.9 with respect to best approximations. This corresponds to [CC16, Corollary 4.4]. In the real case, the characterization works up to multiplication by $1/2$ and by $2$, but the ultrametric nature allows us to obtain the exact equivalence as in Corollary 4.7. The second reason comes from Minkowski’s second theorem (Theorem 2.4). While Minkowski’s second theorem on $\mathbb{R}^{d}$ works with $2^{d}/d!$ and $2^{d}$ multiplications (see [CC16, Equation (M)] and [Cas59]), we have the equality over $\widetilde{\mathcal{K}}^{d}$ as in Theorem 2.4. These two reasons ultimately ensure that the fractal structures we build work well for all $10<\varepsilon<$ 1, unlike [CC16].
@@
Acknowledgments. We would like to thank the anonymous referees for carefully reading this article and providing helpful comments.
@@ 2.1. Basic number theory
Given a non-zero polynomial $f\in\mathcal{R}$, let $f=\alpha P_{1}^{e_{1}}P_{2}^{e_{2}}\cdots P_{t}^{e_{t}}$ be its decomposition to irreducible components, where $\alpha\in\mathbb{F}_{q}^{*}$, $P_{i}$ are distinct monic irreducible polynomials, and $e_{i}$ are non-negative integers. It follows from [Ros13, Propositions 1.7 and 2.4] that
@@ 2.2. Geometry of numbers
[BK23, Lemma 6.2] Let $\Lambda\subseteq\widetilde{\mathcal{K}}^{d}$ be a lattice. Then
@@
$e(X)=\sup\{r>0:\exists\alpha\in\widetilde{\mathcal{K}}^{d}:B(\alpha,r)\cap X=\emptyset\}$
@@
Since $e$ is $\operatorname{SL}_{d}(\mathcal{O})$ invariant, by [Ara24, Theorem 1.24], it suffices to assume that $\Lambda=\mathcal{R}x^{\log_{q}\lambda_{1}(\Lambda)}\mathbf{e}_{1}\oplus\dots\oplus\mathcal{R}x^{\log_{q}\lambda_{d}(\Lambda)}\mathbf{e}_{d}$, where $\mathbf{e}_{1},\dots,\mathbf{e}_{d}$ is the canonical basis for $\widetilde{\mathcal{K}}^{d}$. Define $\mathbf{v}=\sum_{i=1}^{d}x^{\log_{q}\lambda_{i}(\Lambda)-1}\mathbf{e}_{i}$. We claim that
@@
On the other hand, by [Ara24, Theorem 1.25],
$\Lambda+B\left(0,\frac{1}{q}\lambda_{d}(\Lambda)\right)=\widetilde{\mathcal{K}}^{d}.$
@@
Since $e$ is $\operatorname{SL}_{d}(\mathcal{O})$ invariant, then, by [Ara24, Theorem 1.24] , it suffices to assume that $L=\mathcal{R}x^{\log_{q}\lambda_{1}(L)}\mathbf{e}_{1}\oplus\dots\oplus\mathcal{R}x^{\log_{q}\lambda_{d}(L)}\mathbf{e}_{d}$. In this case, Theorem 2.4 implies that
@@ 3. Self-similar coverings
  • for all $\sigma$-admissible sequences $(\alpha_{k})_{k\in\mathbb{N}}$, $\lim_{k\to\infty}\operatorname{diam}B(\alpha_{k})=0$;
@@ 4.1. Best approximation vectors
For each $\boldsymbol{\theta}\in\widetilde{\mathcal{K}}^{d}\setminus\mathcal{K}^{d}$, we associate a sequence $\{\mathbf{u}_{n}\}_{n\in\mathbb{N}}$ of best approximation vectors for $\boldsymbol{\theta}$ with the following properties:
  1. (1)
    $|\mathbf{u}_{1}|=1$;
  2. (2)
    $|\mathbf{u}_{n}|<|\mathbf{u}_{n+1}|$ for all $n\in\mathbb{N}$;
  3. (3)
    there is no best approximation $\mathbf{u}\in Q$ of $\boldsymbol{\theta}$ with $|\mathbf{u}_{n}|<|\mathbf{u}|<|\mathbf{u}_{n+1}|$ for any $n\in\mathbb{N}$.
@@
See [BZ19, KLP23] for the existence of a sequence of best approximation vectors for $\boldsymbol{\theta}\in\widetilde{\mathcal{K}}^{d}\setminus\mathcal{K}^{d}$. We note that multiplying by $\mathbb{F}_{q}^{*}$ does not effect the norm of a vector. Thus, if $\mathbf{u}=(a,b)\in Q$ is a best approximation of $\boldsymbol{\theta}\in\widetilde{\mathcal{K}}^{d}$, then every element of $\mathbb{F}_{q}^{*}\mathbf{u}$ is also a best approximation. Furthermore, $b$ is unique up to $\mathbb{F}_{q}^{*}$-multiples in the following sense. If $\mathbf{v}=(a^{\prime},b^{\prime})\in Q$ satisfies $|b|=|b^{\prime}|$ and $b^{\prime}\notin\mathbb{F}_{q}^{*}b$, then $\mathbf{v}$ is not a best approximation of $\boldsymbol{\theta}$. To see this, assume that $\mathbf{v}$ is a best approximation of $\boldsymbol{\theta}$. Note that there exists some element $\alpha\in\mathbb{F}_{q}^{*}$ such that $0<|b-\alpha b^{\prime}|<|b|$. Since $\mathbf{u}$ and $\alpha\mathbf{v}$ are best approximations of $\boldsymbol{\theta}$ with $|\mathbf{u}|=|\alpha\mathbf{v}|$, we have $A(\boldsymbol{\theta},\mathbf{u})=A(\boldsymbol{\theta},\alpha\mathbf{v})$, hence,
$\|(b-\alpha b^{\prime})\boldsymbol{\theta}-(a-\alpha a^{\prime})\|\leq\max\{A(\boldsymbol{\theta},\mathbf{u}),A(\boldsymbol{\theta},\alpha\mathbf{v})\}=A(\boldsymbol{\theta},\mathbf{u}).$
@@
We note that by Definition 4.1(1), if $\{\mathbf{u}_{n}\}_{n\in\mathbb{N}}$ is a sequence of best approximation vectors, then, the sequence $\{A(\boldsymbol{\theta},\mathbf{u}_{n})\}_{n\in\mathbb{N}}$ is strictly decreasing.
@@
By Remark 4.2, since $b$ is unique up to $\mathbb{F}_{q}^{*}$-multiples, if $b^{\prime}\notin\mathbb{F}_{q}^{*}b$, then we have $A(\boldsymbol{\theta},\mathbf{u})<A(\boldsymbol{\theta},\mathbf{v})$. We note that $b$ is unique as soon as the value $A(\boldsymbol{\theta},\mathbf{u})$ is fixed. Thus, the right hand side of (4.1) is equal to $A(\hat{\mathbf{u}},\mathbf{v})$, that is, $A(\boldsymbol{\theta},\mathbf{v})\leq A(\hat{\mathbf{u}},\mathbf{v})$. Suppose that $b^{\prime}\in\mathbb{F}_{q}^{*}b$ and we may assume that $b^{\prime}=b$ and $A(\boldsymbol{\theta},\mathbf{u})=A(\boldsymbol{\theta},\mathbf{v})$. Since $\mathbf{v}\notin\mathbb{F}_{q}^{*}\mathbf{u}$ and $A(\boldsymbol{\theta},\mathbf{u})\leq 1$, it follows that $A(\hat{\mathbf{u}},\mathbf{v})=\|a-a^{\prime}\|\geq 1\geq A(\boldsymbol{\theta},\mathbf{v})$.
On the other hand, by Lemma 1.5,
@@
Let $\boldsymbol{\theta}\in\widetilde{\mathcal{K}}^{d}\setminus\mathcal{K}^{d}$ and let $\{\mathbf{u}_{n}\}_{n\in\mathbb{N}}$ be a sequence of best approximation vectors for $\boldsymbol{\theta}$. Then, for every $i,j\in\mathbb{N}$, we have
@@ 4.2. Orthogonality and Farey lattice
where $\pi_{\mathbf{u}}(a,b)=b\hat{\mathbf{u}}-a$, where $(a,b)\in\widetilde{\mathcal{K}}^{d}\times\widetilde{\mathcal{K}}$. Note that by Lemma 4.10, $\det(\Lambda_{\mathbf{u}})=|\mathbf{u}|^{-1}$. We define
@@
For $\mathbf{u}\in Q$, let $\Lambda_{\mathbf{u}}^{\prime}$ be the $d-1$ dimensional sublattice of $\Lambda_{\mathbf{u}}$ of minimal covolume, and denote $H_{\mathbf{u}}^{\prime}=\operatorname{span}_{\widetilde{\mathcal{K}}}(\Lambda_{\mathbf{u}}^{\prime})$ and $H_{\mathbf{u}}=\pi_{\mathbf{u}}^{-1}(H_{\mathbf{u}}^{\prime})$. It can be easily checked that $\Lambda_{\mathbf{u}}^{\prime}=\mathrm{span}_{\mathcal{R}}\{\xi_{1},\dots,\xi_{d-1}\}$ and $H_{\mathbf{u}}^{\prime}=\mathrm{span}_{\widetilde{\mathcal{K}}}\{\xi_{1},\dots,\xi_{d-1}\}$.
@@
For $\mathbf{u}\in Q$, we have $\det(\Lambda_{\mathbf{u}})=|\mathbf{u}|^{-1}$ and $\lambda_{1}(\Lambda_{\mathbf{u}})=r(\mathbf{u}).$
@@
In order to prove the first equality, let $\mathbf{u}=(a,b)$ and consider the group homomorphism $\varphi:\Lambda_{\mathbf{u}}\to\mathcal{R}/b\mathcal{R}$ given by
$\varphi\left(\sum_{i=1}^{d}c_{i}\mathbf{e}_{i}+c_{d+1}\hat{\mathbf{u}}\right)=c_{d+1}+b\mathcal{R}.$
Since $\mathrm{gcd}(a,b)=1$, the map $\varphi$ is well-defined. Since $\varphi$ is surjective and $\mathrm{Ker}(\varphi)=\mathcal{R}^{d}$, we have $[\Lambda_{\mathbf{u}}:\mathcal{R}^{d}]=\#\mathcal{R}/b\mathcal{R}=|b|$, where $[\Lambda_{\mathbf{u}}:\mathcal{R}^{d}]$ denotes the index of the subgroup $\mathcal{R}^{d}$ in $\Lambda_{\mathbf{u}}$. Hence, it follows from $\det(\mathcal{R}^{d})=1$ that $\det(\Lambda_{\mathbf{u}})=|b|^{-1}=|\mathbf{u}|^{-1}$.
Now, let us prove the second equality. First, observe that